OpenStudy (anonymous):

3^(x^2+5x+32)=81^(3-x)

OpenStudy (anonymous):

I don't understand @OOOPS

OpenStudy (anonymous):

do you divide (x^2+5x+32) by 3-x?

OpenStudy (anonymous):

well your question wasnt really clear

OpenStudy (mathstudent55):

\(\Large 3^{x^2+5x+32}=81^{3-x}\) \(\Large 3^{x^2+5x+32}=(3^4)^{3-x}\) \(\Large 3^{x^2+5x+32}=3^{12-4x}\) \(\Large x^2+5x+32=12-4x\) Solve the quadratic above.

OpenStudy (anonymous):

@mathstudent55 would i make it into 0?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello!

OpenStudy (misty1212):

\[x^2+5x+32=12-4x\] is what you have to solve right?

OpenStudy (misty1212):

add \(4x\) and subtract \(12\) to get \[x^2+9x+20=0\]

OpenStudy (misty1212):

fortunately for you this one factors so it is easy to solve

OpenStudy (anonymous):

i got (x+5)(x+4) so i then equal it to 0 to find x?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

you can probably do it in your head

OpenStudy (anonymous):

so its -5 and -4

OpenStudy (misty1212):

oui

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (misty1212):

you are welcome dear